OpenStudy (anonymous):

Solve the equation by a method of your choice: x2 + 3x - 4 = 0 If there are two solutions, separate the solutions by a comma.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

is that x^2?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

oh what is it then? x times 2?

OpenStudy (mathstudent55):

Factor the left side.

OpenStudy (mathstudent55):

x2 + 3x - 4 = 0 (x )( ) = 0 Can you factor it?

OpenStudy (anonymous):

2x+3x-4=0 5x-4=0 5x=4 Divide 5 from each side and that is your answer.

OpenStudy (anonymous):

If it is 2x

OpenStudy (mathstudent55):

It's not. It's x^2

OpenStudy (anonymous):

she said it wasn't

OpenStudy (anonymous):

so wat would it be

OpenStudy (mathstudent55):

She may not know what x^2 means. I'm pretty sure she means x^2

OpenStudy (mathstudent55):